OpenStudy (anonymous):

1/2log7x−3log7y−2log5z−log5w

hartnn (hartnn):

properties of log: \(\huge n\log a=log a^n \\ \huge \log(a)-log(b)=log (a/b)\)

hartnn (hartnn):

so \(1/2 log 7x\) can be written as\( log\sqrt{7x} \) ok?

hartnn (hartnn):

3log 7y = log (7y)^3

hartnn (hartnn):

2log 5z = ?